A spinning wheel is slowed down by a brake, giving it a constant angular acceleration of-5.00 rad/s2·orng a 3.40-s time interval, the wheel rotates through 80.4 rad. What is the angular speed of the wheel at the end of the 3.40-s interval? rad/s Need Help? Read it Next
Answer #1

Given is :-

Angular acceleration  \alpha = -5 rad/s^2

time taken  t = 3.40s

Angular displacement = 80.4 radians

Now,

by using the first equation of motion the angular displacement is given by

\theta = w_o t + \frac{1}{2}\alpha t^2

by plugging all the values we get

80.4= w_o (3.40) - \frac{1}{2}(5) (3.40)^2

which gives us initial angular velocity which is

wo = 3215rad/s

Now by using third equation of motion we get

w^2 = w_o^2 + 2 \alpha \theta

by plugging all the values we get

w^2 = (32.15)^2 - 2 (5) (80.4)

which gives us the final angular velocity after the breaking of 3.40 seconds

w-15.15rad/s

Hence the angular speed of the wheel at the end of 3.40 seconds is 15.15 rad/s